Help!!!
Find all real solutions to the following equation.
Square root(symbol) x^2+3x+6=4
Thanks a bunch!!

Answer by wuwei96815(245)

x^2 + 3x + 6 = 4
x^2 + 3x + 6 - 4 = 0
x^2 + 3x + 2 = 0
(x + 1)(x + 2) = 0

Therefore, x = -1, and x = -2 satisfy the equation.